# The Gauss-Markov Adjunction Provides Categorical Semantics of Residuals in Supervised Learning

### v2

@@ 5.2 Diagrams induced from Iterative Updates
are induced from $a_{i+1}:=a_{i}-\eta\nabla_{a_{i}}L(a_{i})$, $L(a_{i}):=\|y-Xa_{i}\|$. The iterative updates of gradient descent converge when the step size coefficient $\eta$ satisfies $0<\eta<2/\lambda_{\max}(XX^{\top})$. These recurrences (26a) and (26b) are related by
@@ B. Iterative Method
Regression can be interpreted as a supervised learning system that learns the optimal parameter $a^{*}$ which enables accurate predictions for the training data $X$, with the value of $y$ serving as the supervisory signal. The objective function is often expressed as $L(a)=\frac{1}{2}\|r\|^{2}=\frac{1}{2}\|y-Xa\|^{2}$. Starting from an initial state $a_{1}$ (arbitrary when ${\rm rank}(X)=m\leq n$ and $a_{1}=0$ when ${\rm rank}(X)\leq n<m$), one can asymptotically approach $a^{*}$ using the iterative method $a_{i+1}=a_{i}-\eta\nabla_{a_{i}}L(a_{i})\;(0<\eta<\eta_{\max}\;;\;i=1,2,\cdots)$, and eventually reach $a^{*}=\lim\limits_{i\to\infty}a_{i}$. Note that $\eta_{\max}:=2/\lambda_{\max}(XX^{\top})$, derived from $\|I-\eta XX^{\top}\|<1$, where $\lambda_{\max}(XX^{\top})=\|X\|^{2}$ denotes the maximum eigenvalue of $XX^{\top}$.

### v3

@@ 5.2 Diagrams induced from Iterative Updates
are induced from $a_{i+1}:=a_{i}-\eta\nabla_{a_{i}}L(a_{i})$, $L(a_{i}):=\|y-Xa_{i}\|$. The iterative updates of gradient descent converge when the step size coefficient $\eta$ satisfies $0<\eta<2/\beta_{\max}(XX^{\top})$. These recurrences (30a) and (30b) are related by
@@ B. Iterative Method
Regression can be interpreted as a supervised learning system that learns the optimal parameter $a^{*}$ which enables accurate predictions for the training data $X$, with the value of $y$ serving as the supervisory signal. The objective function is often expressed as $L(a)=\frac{1}{2}\|r\|^{2}=\frac{1}{2}\|y-Xa\|^{2}$. Starting from an initial state $a_{1}$ (arbitrary when ${\rm rank}(X)=m\leq n$ and $a_{1}=0$ when ${\rm rank}(X)\leq n<m$), one can asymptotically approach $a^{*}$ using the iterative method $a_{i+1}=a_{i}-\eta\nabla_{a_{i}}L(a_{i})\;(0<\eta<\eta_{\max}\;;\;i=1,2,\cdots)$, and eventually reach $a^{*}=\lim\limits_{i\to\infty}a_{i}$. Note that $\eta_{\max}:=2/\beta_{\max}(XX^{\top})$, derived from $\|I-\eta XX^{\top}\|<1$, where $\beta_{\max}(XX^{\top})=\|X\|^{2}$ denotes the maximum eigenvalue of $XX^{\top}$.